OpenStudy (anonymous):

Classify each triangle as acute, right, or obtuse. the lengths are 4,5,7

OpenStudy (harsimran_hs4):



OpenStudy (harsimran_hs4):

cos of an obtuse angle is negative cos A = (b^2 + c^2 - a^2)/2(bc) ||ly for cos B = (a^2 + c^2 - b^2)/2ac etc can you proceed now?